OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -5x – 7y = 25.5 -2x + 10y = -9 I don't understand this elimination method? help?

OpenStudy (mathmagician):

10x+14y=-51 (multiply first equation by -2) -10x+50y=-45 (multiplied by 5) Now add these equations together: 10x+14y-10x+50y=-51-45 64y=-96 y=-96/64=-1.5 10x=-51-14y=-51+21=-30 x=-30/10=-3

OpenStudy (anonymous):

thank you. but how do you know which to eliminate?

OpenStudy (mathmagician):

there is no difference- the result will be the same. For example, you can multiply first equation by 10, second by 7 and add these equations by eliminatinig y. Its up to you